OpenStudy (anonymous):

Extremely quick question! Thank You! Is w=-2-2i in polar form 2sqrt(2)(cos(pi/4)+isin(pi/4))?

OpenStudy (anonymous):

@Ashleyisakitty @satellite73 @jim_thompson5910 @campbell_st @zepdrix @TheSmartOne @Nnesha @Luigi0210 @Compassionate

OpenStudy (anonymous):

I'm worried theta might be wrong

zepdrix (zepdrix):

hmm ya, it can't be pi/4 right? we're in quadrant `three` if both the real and imaginary part are negative.

OpenStudy (anonymous):

-2/-2=1 So I thought that was why. Is there a rule I don't know?

zepdrix (zepdrix):

Well, when you apply your arctangent (im assuming that's what you're doing),\[\Large\rm \arctan\left(\frac{-2}{-2}\right)=\theta\]Yes, you end up with pi/4 as a reference angle. But this really corresponds to two angles in one full rotation. You have to look back at your rectangular form to see where you should be.

zepdrix (zepdrix):

Since they're both negative, we need the NEXT angle, which will be pi/4 + pi

OpenStudy (anonymous):

Oh, I see what you mean.

OpenStudy (anonymous):

5pi/4?

OpenStudy (anonymous):

not to butt in, but drawing a picture is real helpful by which i only mean "plot the point"

zepdrix (zepdrix):

ya 5pi/4 sounds better :)

OpenStudy (anonymous):

Oh, gotcha @satellite73

OpenStudy (anonymous):

it will keep you from doing something silly like taking \[\tan^{-1}(\frac{b}{a})\] which does not really work

OpenStudy (anonymous):

@satellite73 yup, I'll do that next time.

OpenStudy (anonymous):

Thanks for the help everyone :D